OpenStudy (anonymous):

Linear algebra Does anyone know how to convert the vector equation: (1,0) + t(-1,2)=(x,y) into its linear equation, y=mx+b form?

OpenStudy (anonymous):

y=2t x = 1-t t=1-x y=2(1-x) y=-2x+2

OpenStudy (anonymous):

That's right! Could you tell me how you arrived at those equations?

OpenStudy (anonymous):

from (1,0) + t(-1,2)=(x,y)

OpenStudy (anonymous):

got it lol thanks :)